Dave and Rita’s combined age is 46 years. If Dave’s age is 5 years less than twice Rita’s age. How old are Dave and Rita?
Answer by ankor@dixie-net.com(22740)

Dave and Rita’s combined age is 46 years.
d + r = 46
If Dave’s age is 5 years less than twice Rita’s age.
d = 2r - 5
rearrange for elimination
d - 2r = -5
:
d + r = 46
d -2r = -5
---------------subtraction eliminates d, find r
0 + 3r = 51
r = 51/3
r = 17 yrs is Rita's age
then
d + 17 = 46
d = 46 - 17
d = 29 yrs is Dave's age
:
:
Check solution in the statement:
"Dave’s age is 5 years less than twice Rita’s age."
29 = 2(17) - 5
29 = 34 - 5